# Hybrid collective excitations in topological superconductor/ferromagnetic insulator heterostructures

### v3

@@ Appendix B Amplitude response to the magnon
$F_{\omega,\bm{k}}^{a}=\frac{-C^{K}_{12}(\omega,\bm{k})\left(A_{12}^{K*}(-\omega^{*},-\bm{k})-1-B_{12}^{K*}(-\omega^{*},-\bm{k})\right)+C^{K*}_{12}(-\omega^{*},-\bm{k})\left(A_{12}^{K}(\omega,\bm{k})-1-B_{12}^{K}(\omega,\bm{k})\right)}{i\cdot det\left[\hat{M}_{\Delta\Delta}\right]}.$ (104)
@@
$[\delta\hat{g}_{-\omega^{*},-\bm{k}}^{R*}]_{ij}=a_{ij}^{R*}(-\omega^{*},-\bm{k})\delta\Delta_{-\omega,-\bm{k}}^{*}+b_{ij}^{R*}(-\omega^{*},-\bm{k})\delta\Delta_{\omega,k}-c_{ij}^{R*}(-\omega^{*},-\bm{k})(\hat{\bm{k}}\times\delta\bm{h})\cdot\hat{z}$, (105)
@@
$[\delta\hat{g}_{\omega,\bm{k}}^{K}]_{12}=a_{12}^{K}(\omega,\bm{k})\delta\Delta_{\omega,\bm{k}}+b_{12}^{K}(\omega,\bm{k})\delta\Delta_{-\omega,-\bm{k}}^{*}+c_{12}^{K}(\omega,\bm{k})(\hat{\bm{k}}\times\delta\bm{h})\cdot\hat{z}$, (120)
@@
$[\delta\hat{g}_{-\omega,-\bm{k}}^{K}]_{12}^{*}=a_{12}^{K*}(-\omega^{*},-\bm{k})\delta\Delta_{-\omega,-\bm{k}}^{*}+b_{12}^{K*}(-\omega^{*},-\bm{k})\delta\Delta_{\omega,\bm{k}}-c_{12}^{K*}(-\omega^{*},-\bm{k})(\hat{\bm{k}}\times\delta\bm{h}_{-\omega,-\bm{k}})\cdot\hat{z}$, (129)

### v4

@@ Appendix B Amplitude response to the magnon
$F_{\omega_{r},\bm{k}}^{a}=\frac{-C^{K}_{12}(\omega,\bm{k})\left(A_{12}^{K*}(-\omega^{*},-\bm{k})-1-B_{12}^{K*}(-\omega^{*},-\bm{k})\right)+C^{K*}_{12}(-\omega^{*},-\bm{k})\left(A_{12}^{K}(\omega,\bm{k})-1-B_{12}^{K}(\omega,\bm{k})\right)}{i\cdot det\left[\hat{M}_{\Delta\Delta}\right]}.$ (110)
@@
$[\delta\hat{g}_{-\omega_{r},-\bm{k}}^{R*}]_{ij}=a_{ij}^{R*}(-\omega^{*},-\bm{k})\delta\Delta_{-\omega_{r},-\bm{k}}^{*}+b_{ij}^{R*}(-\omega^{*},-\bm{k})\delta\Delta_{\omega_{r},k}-c_{ij}^{R*}(-\omega^{*},-\bm{k})(\hat{\bm{k}}\times\delta\bm{h}_{-\omega_{r},-\bm{k}})\cdot\hat{z}$, (111)
@@
$[\delta\hat{g}_{\omega_{r},\bm{k}}^{K}]_{12}=a_{12}^{K}(\omega,\bm{k})\delta\Delta_{\omega_{r},\bm{k}}+b_{12}^{K}(\omega,\bm{k})\delta\Delta_{-\omega_{r},-\bm{k}}^{*}+c_{12}^{K}(\omega,\bm{k})(\hat{\bm{k}}\times\delta\bm{h}_{\omega_{r},\bm{k}})\cdot\hat{z}$, (126)
@@
$[\delta\hat{g}_{-\omega_{r},-\bm{k}}^{K}]_{12}^{*}=a_{12}^{K*}(-\omega^{*},-\bm{k})\delta\Delta_{-\omega_{r},-\bm{k}}^{*}+b_{12}^{K*}(-\omega^{*},-\bm{k})\delta\Delta_{\omega_{r},\bm{k}}-c_{12}^{K*}(-\omega^{*},-\bm{k})(\hat{\bm{k}}\times\delta\bm{h}_{-\omega_{r},-\bm{k}})\cdot\hat{z}$, (135)